OpenStudy (anonymous):

Find the standard form for the equation of a circle (x-h)^2 + (y-k)^2 = r^2 with a diameter that has endpoints of (-2, 8) and (2, 3).

OpenStudy (ash2326):

In the standard equation \[(x-h)^2+(y-k)^2=r^2\] (h, k) is the center of the circle r= radius we are given the end points of the diameter A(2, -8) and B(2, 3) diameter obviously passes through the center and we can find radius as well if we are given the diameter. @ballerinamd could you find the length of the radius?

OpenStudy (anonymous):

It just occurred to me... 1.) Center point: (h,k) = (0,5.5) [The coordinate (0,5.5) equivalent to the midpoint of the diameter I derived from the endpoints.] Equation of a circle: (x-h)^2+(y-k)^2=r^2 (x-0)^2 + (y-5.5)^2 = r^2 Substitute one of the end-points into the equation to derive the radius: End-point: (2,3) Substitution: (2-0)^2+(3-5.5)^2 = r^2 r^2 = 10.25 r = 3.201 Radius = 3.201 (Thank You!)